Not logged in
Thu Nov 1 01:28:04 2007
First digital lock of OMC
We locked a fiber based NPRO to the suspended OMC tonight using the TPT digital control system. To control the laser frequency, we took the PZT AI output and ran it on a BNC cable down the hallway to the Thorlabs HV box. The Thorlabs is a singled ended unit so we connected the AI positive terminal only and grounded the BNC to the AI shield. We could get a -6 to 1.5 V throw in this method which fed into the 10 k resisotr + 9 V battery at the input of the HV box. The HV out ran to the NPRO PZT fast input.
We derived our error signal from a PDA255 in reflection with a 29.5 MHz PDH lock. The signal feeds into one of the unused Tip/Tilt AA channels and is passed to the PZT LSC drive through the TPT_PDH1 filter bank. In the PZT_LSC filter we put a single pole at 1 Hz which, together with the phase we mentioned the other night (180 degrees at 3 kHz) should allow a 1 kHz-ish loop. In practice, as shown below, we got a 650 Hz UGF with 45 degrees of phase margin and about 6 dB of gain margin.
The Lower figure shows the error point spectrum with 3 settings. REF0 in blue shows lots of gain peaking at 1.5 kHz-ish, just where its expected - the gain was -40. The REF1 has gain of -20 and shows no gain peaking. The current trace in red shows some gain peaking cuz the alignment is better but it also has included a 1^2:20^2 boost which totally crushes the low frequency noise. We should do a better loop sweep after getting the alignment right so we can see how much boost it will really take.
Just for fun, we are leaving it locked overnight and recording the PZT_LSC data for posterity.